# Encoder vs Decoder: Comparative Analysis of Encoder and Decoder Language Models on Multilingual NLU Tasks

### v1

@@ Acknowledgements
Appendix A Appendix
Model ID Decoder Score ($\downarrow$)
AI-Sweden-Models/Llama-3-8B-instruct 1.07
gpt-4-0613 1.28
four-two-labs/lynx-micro 1.29
AI-Sweden-Models/roberta-large-1350k 1.33
meta-llama/Meta-Llama-3-70B 1.34
gpt-4o-2024-05-13 1.35
gpt-4-1106-preview 1.37
KBLab/megatron-bert-large-swedish-cased-165k 1.42
AI-Sweden-Models/bert-large-nordic-pile-1M-steps 1.48
KB/bert-base-swedish-cased 1.48
intfloat/multilingual-e5-large 1.49
ltg/norbert3-large 1.49
gpt-3.5-turbo-0613 1.57
upstage/SOLAR-10.7B-v1.0 1.96
timpal0l/Mistral-7B-v0.1-flashback-v2 2.02
mistralai/Mistral-7B-v0.1 2.17
Table 6: Excerpt of the Swedish ScandEval leaderboard.
Model ID Decoder Score ($\downarrow$)
ltg/norbert3-large 1.12
NbAiLab/nb-bert-large 1.19
ltg/norbert3-base 1.32
AI-Sweden-Models/roberta-large-1350k 1.38
gpt-4-0613 1.49
danish-foundation-models/encoder-large-v1 1.54
google/rembert 1.57
meta-llama/Meta-Llama-3-70B 1.59
gpt-4-1106-preview 1.62
gpt-4o-2024-05-13 1.67
microsoft/mdeberta-v3-base 1.67
gpt-3.5-turbo-0613 2.03
upstage/SOLAR-10.7B-v1.0 2.38
meta-llama/Meta-Llama-3-8B 2.45
mistralai/Mistral-7B-v0.1 2.79
AI-Sweden-Models/Llama-3-8B-instruct 2.82
Table 7: Excerpt of the Norwegian ScandEval leaderboard.
Model ID Decoder Score ($\downarrow$)
vesteinn/FoBERT 1.00
microsoft/mdeberta-v3-base 2.01
vesteinn/ScandiBERT-no-faroese 2.66
gpt-4-1106-preview 2.69
google/rembert 3.35
ltg/norbert3-base 3.36
gpt-4-0613 3.46
mideind/IceBERT-large 3.96
meta-llama/Meta-Llama-3-70B 4.16
gpt-3.5-turbo-0613 4.23
gpt-4o-2024-05-13 4.69
mistralai/Mistral-7B-v0.1 4.84
Table 8: Excerpt of the Faroese ScandEval leaderboard.
Model ID Decoder Score ($\downarrow$)
deepset/gbert-large 1.17
FacebookAI/xlm-roberta-large 1.34
google/rembert 1.34
gpt-4-0613 1.49
intfloat/multilingual-e5-large 1.49
meta-llama/Meta-Llama-3-70B 1.57
gwlms/deberta-base-dewiki-v1 1.61
gpt-4-1106-preview 1.65
upstage/SOLAR-10.7B-v1.0 1.67
microsoft/mdeberta-v3-base 1.74
gpt-4o-2024-05-13 1.82
deepset/gbert-base 1.83
gpt-3.5-turbo-0613 1.95
occiglot/occiglot-7b-de-en-instruct 2.02
meta-llama/Meta-Llama-3-8B 2.02
mistralai/Mistral-7B-v0.1 2.15
Table 9: Excerpt of the German ScandEval leaderboard.
Model ID Decoder Score ($\downarrow$)
intfloat/multilingual-e5-large 1.42
gpt-4-0613 1.69
DTAI-KULeuven/robbert-2022-dutch-base 1.78
pdelobelle/robbert-v2-dutch-base 1.81
meta-llama/Meta-Llama-3-70B 1.84
gpt-4-1106-preview 1.98
intfloat/multilingual-e5-base 2.00
FacebookAI/xlm-roberta-large 2.10
gpt-4o-2024-05-13 2.14
microsoft/mdeberta-v3-base 2.25
gpt-3.5-turbo-0613 2.27
google/rembert 2.29
upstage/SOLAR-10.7B-v1.0 2.33
meta-llama/Meta-Llama-3-8B 2.54
mistralai/Mistral-7B-v0.1 2.77
Table 10: Excerpt of the Dutch ScandEval leaderboard.

### v2

@@ Acknowledgements
Appendix A Prompts
@@
Appendix B Mean Rank Score Algorithm
Algorithm 1 The mean rank score algorithm.
2:Output: The mean rank score of each model
3:$R\leftarrow\textrm{matrix of size }M\times D$
4:for $\hat{d}\in D$ do
6:     for $\hat{m}\in\textrm{sortedModels}$ do
7:         if $\hat{m}$ is the first model then
8:              $\rho\leftarrow 1$
9:              $\textrm{previousScores}\leftarrow S_{\hat{m},\hat{d}}$
10:         else
[…]
12:              if $\textrm{pValue}<0.05$ then
15:                  $\textrm{previousScores}\leftarrow S_{\hat{m},\hat{d}}$
16:              end if
17:         end if
18:         $R_{\hat{m},\hat{d}}\leftarrow\rho$
19:     end for
20:end for
21:$\textrm{output}\leftarrow\nicefrac{{1}}{{D}}\cdot\Sigma_{d}R_{m,d}$
Appendix C Leaderboard Excerpts
Model ID Decoder Score ($\downarrow$)
AI-Sweden-Models/Llama-3-8B-instruct 1.07
gpt-4-0613 1.28
four-two-labs/lynx-micro 1.29
AI-Sweden-Models/roberta-large-1350k 1.33
meta-llama/Meta-Llama-3-70B 1.34
gpt-4o-2024-05-13 1.35
gpt-4-1106-preview 1.37
KBLab/megatron-bert-large-swedish-cased-165k 1.42
AI-Sweden-Models/bert-large-nordic-pile-1M-steps 1.48
KB/bert-base-swedish-cased 1.48
intfloat/multilingual-e5-large 1.49
ltg/norbert3-large 1.49
gpt-3.5-turbo-0613 1.57
upstage/SOLAR-10.7B-v1.0 1.96
timpal0l/Mistral-7B-v0.1-flashback-v2 2.02
mistralai/Mistral-7B-v0.1 2.17
Table 11: Excerpt of the Swedish ScandEval leaderboard.
Model ID Decoder Score ($\downarrow$)
ltg/norbert3-large 1.12
NbAiLab/nb-bert-large 1.19
ltg/norbert3-base 1.32
AI-Sweden-Models/roberta-large-1350k 1.38
gpt-4-0613 1.49
danish-foundation-models/encoder-large-v1 1.54
google/rembert 1.57
meta-llama/Meta-Llama-3-70B 1.59
gpt-4-1106-preview 1.62
gpt-4o-2024-05-13 1.67
microsoft/mdeberta-v3-base 1.67
gpt-3.5-turbo-0613 2.03
upstage/SOLAR-10.7B-v1.0 2.38
meta-llama/Meta-Llama-3-8B 2.45
mistralai/Mistral-7B-v0.1 2.79
AI-Sweden-Models/Llama-3-8B-instruct 2.82
Table 12: Excerpt of the Norwegian ScandEval leaderboard.
Model ID Decoder Score ($\downarrow$)
vesteinn/FoBERT 1.00
microsoft/mdeberta-v3-base 2.01
vesteinn/ScandiBERT-no-faroese 2.66
gpt-4-1106-preview 2.69
google/rembert 3.35
ltg/norbert3-base 3.36
gpt-4-0613 3.46
mideind/IceBERT-large 3.96
meta-llama/Meta-Llama-3-70B 4.16
gpt-3.5-turbo-0613 4.23
gpt-4o-2024-05-13 4.69
mistralai/Mistral-7B-v0.1 4.84
Table 13: Excerpt of the Faroese ScandEval leaderboard.
Model ID Decoder Score ($\downarrow$)
deepset/gbert-large 1.17
FacebookAI/xlm-roberta-large 1.34
google/rembert 1.34
gpt-4-0613 1.49
intfloat/multilingual-e5-large 1.49
meta-llama/Meta-Llama-3-70B 1.57
gwlms/deberta-base-dewiki-v1 1.61
gpt-4-1106-preview 1.65
upstage/SOLAR-10.7B-v1.0 1.67
microsoft/mdeberta-v3-base 1.74
gpt-4o-2024-05-13 1.82
deepset/gbert-base 1.83
gpt-3.5-turbo-0613 1.95
occiglot/occiglot-7b-de-en-instruct 2.02
meta-llama/Meta-Llama-3-8B 2.02
mistralai/Mistral-7B-v0.1 2.15
Table 14: Excerpt of the German ScandEval leaderboard.
Model ID Decoder Score ($\downarrow$)
intfloat/multilingual-e5-large 1.42
gpt-4-0613 1.69
DTAI-KULeuven/robbert-2022-dutch-base 1.78
pdelobelle/robbert-v2-dutch-base 1.81
meta-llama/Meta-Llama-3-70B 1.84
gpt-4-1106-preview 1.98
intfloat/multilingual-e5-base 2.00
FacebookAI/xlm-roberta-large 2.10
gpt-4o-2024-05-13 2.14
microsoft/mdeberta-v3-base 2.25
gpt-3.5-turbo-0613 2.27
google/rembert 2.29
upstage/SOLAR-10.7B-v1.0 2.33
meta-llama/Meta-Llama-3-8B 2.54
mistralai/Mistral-7B-v0.1 2.77
Table 15: Excerpt of the Dutch ScandEval leaderboard.